OpenStudy (anonymous):

Please someone check this for me: Simplify. -(10)^-1 = -10 (-2)^-5 = 1/32 4^8 x 4^3 = 4^11 (-4)^-6 x (-4)^7 = -4 or 1/4 (3^7)^2 = 3^14 4c^-1 x 3c^10 = 12c^9 (-h^4)^5 = -h^20

OpenStudy (bahrom7893):

-(10)^-1 = -10 is not true

OpenStudy (anonymous):

just that one?

OpenStudy (bahrom7893):

(-2)^-5 = 1/32 is not true either

OpenStudy (anonymous):

and would it be positive 10?

OpenStudy (bahrom7893):

no it would be -1/10

OpenStudy (bahrom7893):

(-2)^-5=-1/32

OpenStudy (anonymous):

okay, are you sure because this is going to be on my test and i don't wanna fail lol :)

OpenStudy (anonymous):

and thank you!

OpenStudy (bahrom7893):

yes, im sure. wait let me check them all

OpenStudy (anonymous):

okay thanks! =)

OpenStudy (anonymous):

-4 is not the same as 1/4

OpenStudy (anonymous):

so just -4? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thanks.

OpenStudy (bahrom7893):

luwcey is going to explain why.

OpenStudy (anonymous):

i have few other problems i don't know how to do, if you have time of course.

OpenStudy (anonymous):

It's really hard to explain & yes justinee I can help you

OpenStudy (anonymous):

(-5a^3b^6)^3(a^4b^2)^7

OpenStudy (anonymous):

(-15a^9b^18)(a^28b^14) = -15a^37b^22

OpenStudy (anonymous):

i thought -5^3 = -125

OpenStudy (anonymous):

you're right! my bad sorry!

OpenStudy (anonymous):

oh okay!:) thank you! & can you help me with 2 more problems if you can.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):